# Apparently ultra-long period radio sources from self-lensed pulsar-black hole binaries

### v2

@@ Abstract
Pulsar-black hole (BH) close binary systems, which have not been found yet, are unique laboratories for testing theories of gravity and understanding the formation channels of gravitational-wave sources. We study the self-gravitational lensing effect in a pulsar-BH system on the pulsar’s emission. Because this effect occurs once per orbital period for almost edge-on binaries, we find that it could generate apparently ultra-long period (minutes to hours) radio signals when the intrinsic pulsar signal is too weak to detect. Each of such lensed signals, or ’pulse’, is composed of a number of amplified intrinsic pulsar pulses. We estimate that a radio telescope with a sensitivity of $10\,\rm mJy$ could detect $\sim$ a few systems that emit such signals in our galaxy. The model is applied to three recently found puzzling long-period radio sources: GLEAM-X J1627, PSR J0901-4046, and GPM J1839-10. To explain their observed signal durations and periods, the masses of their lensing components would be $\sim 10^{4}\,\rm M_{\odot}$, $\sim 4\,\rm M_{\odot}$ and $10^{3-6}\,\rm M_{\odot}$, respectively, with their binary coalescence times ranging from a few tens to thousands of years. However, the implied merger rates (as high as $\sim 10^{3-4}\,\rm Myr^{-1}$ per galaxy) and the large period decay rates ($>10^{-8}\,\rm s\,s^{-1}$) tend to disfavour this self-lensing scenario for these three sources. Finally, for a binary containing a millisecond pulsar and a stellar-mass BH, the Shapiro delay effect would cause a $\geq 10\%$ variation of the profile width for the sub-pulses in such lensed signals.
@@ 4.1 Puzzling radio sources
The above case applications of our self-lensing scenario to GLEAM-X J1627 and GPM J1839-10 suggest they are intermediate mass ratio inspirals (IMRIs), consisting of an NS and an intermediate-mass black hole (IMBH). One of the most promising environments to embed IMBHs is the centres of star clusters. Since the formation of an IMBH is related to the numerous low-mass compact objects that orbit around it, star clusters are also ideal places to find IMRIs.
@@ 6 Discussion and Conclusion
We estimate that, in the Milky Way, there would be at least $12$ NS-BH binaries that can reach a threshold amplification of $10$. Observationally, there would be $\sim 3$ pulsar-BH binaries with self-lensing signals observable for a telescope with a sensitivity of $10\,mJy$, and $\sim 2$ of them would exhibit ultra-long period radio signals. Because of the assumption that all pulsars have a uniform radio luminosity, the actual number of observable sources may be even lower due to the existence of numerous candidates with luminosity lower than the value selected. However, telescopes with lower sensitivities ($\sim\rm\mu Jy$ for long period signals), such as FAST (Han et al., 2021) and SKA (Braun et al., 2019), hold the potential to identify more sources that contain a low luminosity pulsar, increasing the possibility of detecting such systems.
Three recently found long-period radio sources, GLEAM-X J1627 (Hurley-Walker et al., 2022), PSR J0901-4046 (Caleb et al., 2022) and GPM J1839-10 (Hurley-Walker et al., 2023) are puzzling about their physical origins. We interpret them as self-lensed pulsar-BH binaries. Our modeling of them suggests BH masses of $\sim 10^{4}\,\rm M_{\odot}$ and $\sim 4\,\rm M_{\odot}$ for the former two sources, with coalescence times of $\sim 6\times 10^{2}\,\rm yrs$ and $\sim 10^{2}\,\rm yrs$, respectively, basing on their inferred orbital separations. Given the substantial variation of the signal duration, the BH mass of the third source, GBM J1839-10 is estimated to range between $10^{3}$ and $10^{6}\,\rm M_{\odot}$, with a coalescence time spanning $\sim 30$ to $\sim 3\times 10^{3}\,\rm yrs$, implying at least 30 years more of activity, which could be verified by future observation.
Our model predicts negative, large values ($\sim 10^{-8}$ s s-1) of the period derivative for the three sources. However, their observed $\dot{P}$’s are all much smaller ($\lesssim 10^{-9}$ s s-1) and even favour positive values (slowing down) for the source PSR J0901-4046. This discrepancy suggests that other physical processes might be involved. For instance, the spin angular momentum of the pulsar might be transferred to the orbital one, via tidal or some sort of magnetic braking effect, thus reversing the orbital period decay.
@@
This work is supported by National Natural Science Foundation of China (grants 12073091, 12261141691 and 12393814). We thank the anonymous reviewer whose comments have helped improve the quality of the manuscript. We also thank the comments from Baixuan Tan and Long Ji.

### v3

# Apparently ultra-long period radio signals from self-lensed pulsar-black hole binaries

@@ Abstract
Pulsar-black hole (BH) close binary systems, which have not been found yet, are unique laboratories for testing theories of gravity and understanding the formation channels of gravitational-wave sources. We study the self-gravitational lensing effect in a pulsar-BH system on the pulsar’s emission. Because this effect occurs once per orbital period for almost edge-on binaries, we find that it could generate apparently ultra-long period (minutes to hours) radio signals when the intrinsic pulsar signal is too weak to detect. Each of such lensed signals, or ’pulse’, is composed of a number of amplified intrinsic pulsar pulses. We estimate that a radio telescope with a sensitivity of $10\,\rm mJy$ could detect $\sim$ a few systems that emit such signals in our galaxy. The model is applied to three recently found puzzling long-period radio sources: GLEAM-X J1627, PSR J0901-4046, and GPM J1839-10. To explain their observed signal durations and periods, the masses of their lensing components are estimated as $\sim 10^{4}\,\rm M_{\odot}$, $\sim 4\,\rm M_{\odot}$ and $10^{3-6}\,\rm M_{\odot}$, respectively, with their binary coalescence times ranging from a few tens to thousands of years. However, the implied merger rates (as high as $\sim 10^{3-4}\,\rm Myr^{-1}$ per galaxy) and the large period decay rates ($>10^{-8}\,\rm s\,s^{-1}$) tend to disfavour this self-lensing scenario for these three sources. Despite this, our work still provides observational characteristics for self-lensed pulsar-BH binaries, which could help the detection of related sources in the future. Finally, for a binary containing a millisecond pulsar and a stellar-mass BH, the Shapiro delay effect would cause a $\geq 10\%$ variation of the profile width for the sub-pulses in such lensed signals.
@@ 4.1 Puzzling radio sources
Assuming the self-lensing scenario we describe above, GLEAM-X J1627 and GPM J183910 are interpreted as intermediate-mass ratio inspirals (IMRIs), consisting of an NS and an intermediate-mass black hole (IMBH). One of the most promising environments to embed IMBHs is the centres of star clusters. Since the formation of an IMBH is related to the numerous low-mass compact objects that orbit around it, star clusters are also ideal places to find IMRIs.
@@ 6 Discussion and Conclusion
We estimate that, in the Milky Way, there would be at least $12$ NS-BH binaries that can reach a threshold amplification of $10$. Observationally, there would be $\sim 3$ pulsar-BH binaries with self-lensing signals observable for a telescope with a sensitivity of $10\,\rm mJy$, and $\sim 2$ of them would exhibit ultra-long period radio signals. Because of the assumption that all pulsars have a uniform radio luminosity, the actual number of observable sources may be even lower due to the existence of numerous candidates with luminosity lower than the value selected. However, telescopes with lower sensitivities ($\sim\rm\mu Jy$ for long period signals), such as FAST (Han et al., 2021) and SKA (Braun et al., 2019), hold the potential to identify more sources that contain a low luminosity pulsar, increasing the possibility of detecting such systems.
Three recently found long-period radio sources, GLEAM-X J1627 (Hurley-Walker et al., 2022), PSR J0901-4046 (Caleb et al., 2022) and GPM J1839-10 (Hurley-Walker et al., 2023) are puzzling about their physical origins. We apply our model to these sources. Assuming they are self-lensed pulsar-BH binaries, our modelling suggests BH masses of $\sim 10^{4}\,\rm M_{\odot}$ and $\sim 4\,\rm M_{\odot}$ for the former two sources, with coalescence times of $\sim 6\times 10^{2}\,\rm yrs$ and $\sim 10^{2}\,\rm yrs$, respectively, basing on their inferred orbital separations. Given the substantial variation of the signal duration, the BH mass of the third source, GBM J1839-10 is estimated to range between $10^{3}$ and $10^{6}\,\rm M_{\odot}$, with a coalescence time spanning $\sim 30$ to $\sim 3\times 10^{3}\,\rm yrs$, implying at least 30 years more of activity, which could be verified by future observation.
Our model predicts negative, large values ($\sim 10^{-8}$ s s-1) of the period derivative for the three sources. However, their observed $\dot{P}$’s are all much smaller ($\lesssim 10^{-9}$ s s-1) and even favour positive values (slowing down) for the source PSR J0901-4046. Under the self-lensing scenario, this discrepancy could result from other physical processes. For instance, the spin angular momentum of the pulsar might be transferred to the orbital one, via tidal or some sort of magnetic braking effect, thus reversing the orbital period decay.
@@
Although the implied high merger rate and the observed period derivative of these three sources tend to disfavour the self-lensing model for them, our work still provides distinguishable observation appearances (§2) and the number prediction (§3) for those self-lensed pulsar-BH binaries. It still holds the potential to detect the self-lensed pulsar-BH binaries in the future, especially with more sensitive facilities.
@@
This work is supported by National Natural Science Foundation of China (grants 12073091, 12261141691 and 12393814). We thank the anonymous reviewer whose comments have helped improve the quality of the manuscript. We also thank the comments from Tam, Pak-Hin Thomas and Long Ji.